Circular No. 6130

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

NOVA CIRCINI 1995 W. Liller, Instituto Isaac Newton, Viña del Mar, reports his discovery

on a single PROBLICOM photograph (85-mm lens, orange filter, Kodak TP film) taken Jan. 27.328 UT of a stellar object of m_{v} ∼ 7.2 at the position α = 14^{h}44^{m}58^{s}, δ = −63^{o}54′.0 (equinox 2000.0). Nothing brighter than mag 12.0 appears at this position on a film taken Jan. 12.336 UT. The sharpness of the image suggests that it is not a red variable.

S. Benetti, European Southern Observatory (ESO); and E. Cappellaro, Padua Observatory, report that inspection of a fully-reduced CCD spectrum (range 390–900 nm, resolution 1.3 nm) obtained on Jan. 30.38 UT with the ESO New Technology Telescope (+ EMMI) confirms that this object is a nova. The spectrum is dominated by intense emission lines of hydrogen, Fe II, and O I (777.4 and 844.6 nm). The FWHM of Hα, Hβ, and Hγ, corrected for the instrumental resolution, is ∼ 1500 km/s.

A. C. Gilmore, Mount John University Observatory, reports that N Cir 1995 shows nova-like colors in photometry obtained at Mt. John with the 0.6-m f/13 reflector (conditions were good but the region low): Jan. 29.449 UT, V = 7.91 ± 0.02, U −B = −0.44 ± 0.50, B −V = +0.38 ± 0.02, V −R = +1.50 ± 0.02, V − I = +1.71 ± 0.02 (airmass 2.14). Comparison was with two stars in Cousins standard region F2 and nearby HR 5539. The true U − B is likely to be more negative than obtained here.

Visual magnitude estimates from Australia reported by P. Nelson: Jan. 28.526 UT, 7.4 (R. Stubbings); 29.464, 8.1 (Nelson).

EF PEGASI Further visual magnitude estimates by P. Schmeer, Bischmisheim, Ger-

many (cf. IAUC 6125): Jan. 12.73 UT, 11.3; 13.72, 11.3; 13.74, 11.3; 16.741, 11.6; 17.720, 11.8; 19.747, 12.0.

COMET 19P/BORRELLY Total visual magnitude estimates: 1994 Dec. 16.20 UT, 9.0 (A. Kósa-

Kiss, Salonta, Romania, 0.16-m reflector); 29.26, 8.9 (A. Hale, near Chaparral, NM, 10×50 binoculars); 1995 Jan. 1.18, 9.1 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 6.20, 9.0 (Spratt); 14.16, 9.6 (Spratt); 22.16, 10.0 (Spratt); 26.19, 10.2 (Spratt).

1995 January 30 Daniel W. E. Green